OpenStudy (anonymous):

A town of 10,000 people has been growing by 300 people each year A second town of 30,000 has been shrinking by 200 people each year.  When will the two towns have the same population?

OpenStudy (anonymous):

10000+300x=30000-200x 300x+200x=30000-10000 500x=20000 x=40

OpenStudy (anonymous):

hey thanks!